Solve for the variable. Show all work please. Please also include a check. If the solution is undefined please say so. Thanks :)
(1-y)/(1+y)=(2)/(3)

Found 2 solutions by stanbon, lynnlo:
Answer by stanbon(75887)

(1-y)/(1+y)=(2)/(3)
----
Cross-multiply:
3(1-y) = 2(1+y)
-----
3-3y = 2 + 2y
----
5y = 1
y = 1/5
------------------
Cheers,
Stan H.

Answer by lynnlo(4176)

1-y=2
===========
1+y= 3
2(1+y)=3(1-y)
2+2y=3-3y
5y=1
y=1/5